OpenStudy (anonymous):

Find the product:(x-2)(x^2+2x+4) A.x^3-4x+4 B.x^2+3x+2 C.-2x^3-4x^2-8 D.x^3-8

OpenStudy (anonymous):

difference of cubes

OpenStudy (aroub):

yes which is D.

OpenStudy (anonymous):

(a-b)(a^2+ab+b^2)=a^3-b^3 a=x b=2